OpenStudy (anonymous):

Let E be the point where the line l intersects the plane p. Find, in the scalar parametric equation for the line, the value of the parameter which corresponds to the point E and hence find the co-ordinates of this point.

OpenStudy (anonymous):

"scalar parametric"? In 3D, there is no scalar equation of a line because no unique normal. Do you just mean parametric (x=x_0+at, etc)?

OpenStudy (anonymous):

yes that's the one. Also i forgot to include the information i already worked out for the questions before this one which i should probably specify! equation of the plane -x-y+4z=3 equation of normal vector to the plane n=-i-j+4k scalar parametric equations of a line perpendicular to the plane which passes through the point (-1,1,1) ... (x=-1-t),(y=1-t),(z=1+4t)

OpenStudy (anonymous):

the line (x=-1-t), (y=1-t), (z=1+4t) is the line i'm talking about in the question

OpenStudy (dumbcow):

given a line and a plane, solve for t that will yield point of intersection line: \[x = x_0+\alpha t\] \[y=y_0+\beta t\] \[z =z_0 +\gamma t\] plane: \[ax+by+cz+d = 0\] by plugging in line equations into plane equation you can find point that is both in plane and on line then solving for t \[t = \frac{-(ax_0+by_0+cz_0+d)}{a \alpha +b \beta +c \gamma}\]

OpenStudy (anonymous):

thankyou! but what does aα+bβ+cγ correspond to?

OpenStudy (dumbcow):

they are all scalar coefficients a,b,c refer to plane coefficients alpha,beta,gamma refer to line coefficients

OpenStudy (dumbcow):

oh in your case, since line is perpendicular a = alpha b = beta c = gamma

OpenStudy (anonymous):

that is fantastic and so helpful, thank you! you've saved me! would you mind helping with another question if you've got time?

OpenStudy (dumbcow):

sure i can take a look at it

OpenStudy (anonymous):

line 1: \[x=1+2r\] \[y=1+2r\] \[z=-3+r\] line 2: \[x=-2\] \[y=-s\] \[z=2+s\] Find the unit vector nˆ with negative i component which is perpendicular to both line1 and line2 i only know how to take the cross product of ordinary vectors to get a unit vector, i'm not sure what to do with parametric equations?

OpenStudy (dumbcow):

ok for parametric equations , look at directional vectors which are the "t" coefficients if the dot product of 2 vectors equals 0 then they are perpendicular line 1: 2i+2j+k line 2: 0i-j+k let the unit vector be , ai +bj +ck where a <0 and a^2 +b^2 +c^2 = 1 (unit vector) now apply dot product 2a+2b+c = 0 0 -b +c = 0 find (a,b,c) that satisfies all conditions

OpenStudy (dumbcow):

i have to get off OS now...here is the solution \[<a,b,c> = <-\frac{3\sqrt{17}}{17},\frac{2\sqrt{17}}{17},\frac{2\sqrt{17}}{17}>\]

OpenStudy (anonymous):

Thanks so much I really was stuck and you have helped a great deal I am in your debt !

OpenStudy (dumbcow):

your very welcome